OpenStudy (bortain36):

Solve (Square root of 6)^8 x = 216^x−3.

OpenStudy (bortain36):

I'm sorry, I don't know how to put a square root in this equation

jhonyy9 (jhonyy9):

so this wan being ((sqrt6)^8)x or x is exponent too ? and on the other side what is exponent just x or x-3 ?

OpenStudy (bortain36):

Yes, x is apart of the exponent, actually I will send photo

jhonyy9 (jhonyy9):

ok and on the right side ?

OpenStudy (bortain36):

Here is photo

jhonyy9 (jhonyy9):

ok now is more clearly so there are (sqrt6)^8x = 216^(x-3)

jhonyy9 (jhonyy9):

(sqrt6)^8x = (6)^((1/2)*8x) = 6^(8x/2) =6^(4x) and on the other side do you can rewriting the 216 like a power of 6 ?

jhonyy9 (jhonyy9):

in this way 6^2 = 36 so 6^3 = 36*6 = ?

jhonyy9 (jhonyy9):

use a calculator please collaborate

OpenStudy (bortain36):

Yes I see

jhonyy9 (jhonyy9):

so than how will be 216 = 6^(?)

OpenStudy (bortain36):

216

OpenStudy (bortain36):

4

OpenStudy (bortain36):

I mean 3

jhonyy9 (jhonyy9):

yes 3 is right

jhonyy9 (jhonyy9):

so than you can writing

jhonyy9 (jhonyy9):

6^(4x) = 6^3 yes ?

OpenStudy (bortain36):

yes

jhonyy9 (jhonyy9):

so than you see the 6 base on the other sides are the same so how you think when will be this sides equal ?

OpenStudy (bortain36):

Yes, they will be equal

jhonyy9 (jhonyy9):

ok but when ? how need being these exponents too same with the base of 6 ?

OpenStudy (bortain36):

We divide yes?

OpenStudy (bortain36):

Now we have 4x=3

jhonyy9 (jhonyy9):

so looke please 6^x =6^y just in case than x=y do you understand me now ?

jhonyy9 (jhonyy9):

yes exactly how you wrote above

jhonyy9 (jhonyy9):

4x =3 than x = ?

OpenStudy (bortain36):

3/4

jhonyy9 (jhonyy9):

but there are 216^(x-3) yes ? what we forgot it

jhonyy9 (jhonyy9):

so than 216 = 6^3 so there will be 6^3(x-3) yes ?

jhonyy9 (jhonyy9):

so in this case will result 4x = 3(x-3) do you can solving it now for x ?

jhonyy9 (jhonyy9):

are you here please ?

OpenStudy (bortain36):

yes

OpenStudy (bortain36):

Internet is slow

jhonyy9 (jhonyy9):

ok wait

OpenStudy (bortain36):

4x=3x-9 x=-9

jhonyy9 (jhonyy9):

yes you are right but do you understand it now sure clearly ?

OpenStudy (bortain36):

Yes, thank you

OpenStudy (bortain36):

I appreciate this very much

jhonyy9 (jhonyy9):

welcome was my pleasure bye bye good luck

OpenStudy (bortain36):

Bye